OpenStudy (anonymous):

find the x-intercepts of the parabola with vertex(6,27) and y-intercept (0,-81). write anwser in form (x1,y1)(x2,y2). if necessary ,round to the nearest hundredth . :D PLEASE HELP ME ! <3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

since the vertex is \[(6,27)\] you know it looks like \[y=a(x-6)^2+27\] so now you need a

OpenStudy (anonymous):

if x = 0, you know y = -81 so replace x by 0 and set y = -81 to get \[-81=a(0-6)^2+27\] \[-81=36a+27\] \[-108=36a\] \[a=\frac{-108}{36}=-3\]

OpenStudy (anonymous):

so your equation is \[y=-3(x+6)^2+27\] or if you prefer \[y-27=-3(x+6)^2\]

OpenStudy (anonymous):

oh ok i understand thanks but love what is it in x1,y1 form x2,y2 form ?

OpenStudy (anonymous):

since the parabola is shifted to the right, shouldn't the equation be -3(x-6)^2 + 27

OpenStudy (anonymous):

the x intercepts are 3 and 9...

OpenStudy (anonymous):

so you have to write them in as \[(3,0)\] and \[(9,0)\]